In a retail store, the owner earns 1/4 more than the manager earns. The manager is earning what percentage of the owner's salary?

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)
Answer by richwmiller(17219)

o=1.25*m
o=5/4*m
4/5*o=m
4/5 =80% owners salary not 75%